# FFAA: Multimodal Large Language Model based Explainable Open-World Face Forgery Analysis Assistant

### v1

@@ Introduction
MIDS consists of three components: Feature Encoding, Image-Answer Cross Fusion Module and Classification Layer, as depicted in Fig. 4b. We first utilize the fine-tuned MLLM to extract answers from unused samples in the Multi-attack dataset for training. Specifically, for a given image $\bm{X}_{v}$, a non-hypothetical prompt is randomly selected to query the MLLM, yielding $\bm{X}_{a}^{a}$. Based on the determination in $\bm{X}_{a}^{a}$, two hypothetical prompts assuming the same or opposite determination are used to query again, yielding the positive answer $\bm{X}_{a}^{p}$ and the negative answer $\bm{X}_{a}^{n}$ respectively. Finally, based on the authenticity $y_{v}\in\{0,1\}$ of $\bm{X}_{v}$ and the determination result $y_{a}\in\{0,1\}$ of $\bm{X}_{a}$, we assign a label $y$ for each $(\bm{X}_{v},\bm{X}_{a})$ as:
@@
For each given answer $\bm{X}_{a}$, to prevent the model from relying solely on the final classification result while ignoring the intermediate analysis process, we mask the ’Analysis result’ of $\bm{X}_{a}$. Then, we employ the pre-trained T5-Encoder (Raffel et al. 2020) with frozen weights for good linguistic consistency to obtain the text feature $\bm{F}_{t}$. For the image $\bm{X}_{v}$, we employ the pre-trained CLIP-ViT-L/14 (Radford et al. 2021) as the visual feature encoder. To maintain basic image understanding and enhance the perception of local and global features related to face authenticity, the weights of the last two layers are unfrozen. Then, we employ trainable projection matrices $(\bm{W}_{vg},\bm{W}_{vl})$ to project the output features of the last two layers into the same dimensional space as $\bm{F}_{t}$, resulting in $\bm{F}_{vl}$ and $\bm{F}_{vg}$.
Image-Answer Cross Fusion.
@@
To extract the corresponding visual features based on the answer and to identify matching textual features from the visual features, we employ dual cross-attention for comprehensive fusion across all dimensions. Specifically, it can be expressed as:
$\bm{F}_{v\to t}=LayerNorm(softmax(\frac{\bm{Q}_{v}\bm{K}_{t}^{T}}{\sqrt{d}})\bm{V}_{t}),$ (4)
$\bm{F}_{t\to v}=LayerNorm(softmax(\frac{\bm{Q}_{t}\bm{K}_{v}^{T}}{\sqrt{d}})\bm{V}_{v}),$
@@
where $d$ is the dimension of these features, $\{\bm{Q}_{t},\bm{K}_{t},\bm{V}_{t}\}=\mathcal{F}_{qkv}^{t}(\bm{F}_{t})$, $\{\bm{Q}_{v},\bm{K}_{v},\bm{V}_{v}\}=\mathcal{F}_{qkv}^{v}(\bm{F}_{v})$. We perform the operations of Eq. 4 with $(\bm{F}_{t},\bm{F}_{vl})$ and $(\bm{F}_{t},\bm{F}_{vg})$ respectively, to obtain $\bm{F}_{t\to vl}$, $\bm{F}_{vl\to t}$, $\bm{F}_{t\to vg}$ and $\bm{F}_{vg\to t}$. Subsequently, we employ learnable vectors $\bm{f}^{0}$ to obtain the corresponding aggregated feature based on cross-attention:
$\bm{f}^{r}=LayerNorm(softmax(\frac{\bm{Q}_{f}\bm{K}_{F}^{T}}{\sqrt{d}})\bm{V}_{F}),$ (5)
$\bm{Q}_{f}=\bm{W}_{f}^{Q}\bm{f}^{r-1},~{}\bm{K}_{F}=\bm{W}_{F}^{K}\bm{F},~{}\bm{V}_{F}=\bm{W}_{F}^{V}\bm{F},~{}$
@@
where $r$ denotes the $r$-th iteration. Through the operations of Eq. 5, we obtain text feature-enhanced visual domain features $(\bm{f}_{t\to vl}^{r},\bm{f}_{t\to vg}^{r})$ and image feature-enhanced textual domain features $(\bm{f}_{v\to tl}^{r},\bm{f}_{v\to tg}^{r})$ for each $(\bm{X}_{v},\bm{X}_{a})$.
Classification and Loss Function.
@@
To select the best answer, the model must possess a certain ability to discern the authenticity of faces and answers. Therefore, we set the classification labels as the four-class categorization shown in Eq. 3. The [CLS] tokens of $\bm{F}_{vl}$ and $\bm{F}_{vg}$, containing enriched image category information, are passed through projection matrices $(\bm{W}_{vg}^{cls},\bm{W}_{vl}^{cls})$ to obtain $(\bm{f}_{vl}^{cls},\bm{f}_{vg}^{cls})$. Then, we concatenate $(\bm{f}_{vg}^{cls},\bm{f}_{t\to vg}^{r},\bm{f}_{v\to tg}^{r},\bm{f}_{vl}^{cls},\bm{f}_{t\to vl}^{r},\bm{f}_{v\to tl}^{r})$ to obtain $\bm{H}_{f}$, which are further enhanced by $M$ iterations of self-attention with residual connections (He et al. 2016), resulting in $\bm{H}^{\prime}_{f}$. Finally, $\bm{H}^{\prime}_{f}$ is passed through $\bm{W}_{c}$ and a softmax to obtain $\bm{m}\in\mathbb{R}^{4}$. The loss is computed as follows:
@@
Source Deeperforensics-1.0 DeepFakeDetection DFDC
@@
Source PGGAN, Celeb-A WhichFaceIsReal MultiFFDI
@@
Figure 5: Qualitative example. The first row presents a forged face image sourced from DFDC, the prompt, and the response from LLaVA-v1.6-34B. The second and third rows display the responses from GPT-4o and FFAA, respectively. Judgment results are indicated in parentheses as (’Real’, ’Fake’, ’Refuse to judge’), with green for correct judgments and red for incorrect ones.
[…]
Xception 61.1 / 60.2 / 70.8 78.0 / 87.8 / 94.4 45.7 / 74.8 / 91.9 69.4 / 96.5 / 93.6 91.7 / 98.5 / 98.7 79.3 / 86.5 / 88.1 70.9 / 75.2 / 83.7 96.8 / 99.4 / 99.8 16.3
RECCE 39.9 / 28.7 / 46.0 72.2 / 82.8 / 91.7 41.5 / 73.7 / 92.4 93.6 / 98.0 / 96.6 91.6 / 97.4 / 97.4 75.9 / 83.6 / 87.3 69.1 / 74.0 / 81.7 98.1 / 99.6 / 99.9 22.4
Implicit* 46.7 / 46.5 / 50.3 77.5 / 87.7 / 94.1 53.7 / 70.8 / 91.5 56.6 / 67.0 / 45.3 49.6 / 51.1 / 51.7 67.8 / 78.1 / 79.0 58.7 / 63.3 / 69.7 67.8 / 70.2 / 88.4 10.4
FedForgery* 75.8 / 83.5 / 78.9 76.6 / 88.1 / 73.5 38.6 / 58.2 / 25.0 68.5 / 55.9 / 83.4 48.9 / 47.8 / 51.2 52.5 / 56.8 / 58.2 60.1 / 67.8 / 62.8 51.0 / 64.6 / 33.1 13.7
Exposing 64.2 / 75.5 / 80.7 76.4 / 86.2 / 93.3 48.6 / 77.2 / 93.0 62.3 / 94.8 / 87.1 90.0 / 98.0 / 98.1 75.2 / 83.4 / 84.1 69.5 / 76.8 / 82.6 94.5 / 98.8 / 99.6 14.9
NPR 47.5 / 44.2 / 52.5 67.2 / 74.5 / 85.9 40.1 / 66.4 / 88.6 67.2 / 80.1 / 56.2 91.6 / 99.3 / 99.3 71.4 / 79.4 / 80.4 64.2 / 72.5 / 76.1 97.1 / 99.7 / 99.9 19.3
CLIPping 90.4 / 97.6 / 98.0 77.8 / 86.6 / 94.0 64.0 / 71.9 / 92.3 46.0 / 97.1 / 92.4 91.5 / 98.0 / 98.1 55.6 / 96.4 / 96.3 70.9 / 80.9 / 84.6 87.4 / 94.2 / 98.0 16.9
FFAA (Ours) 94.2 / 98.8 / 98.9 82.7 / 92.0 / 96.4 65.0 / 74.0 / 92.3 91.8 / 99.7 / 99.3 98.2 / 99.8 / 99.9 87.1 / 98.6 / 98.9 86.5 / 94.4 / 95.5 86.8 / 96.7 / 98.9 10.0
Table 2: Comparison results (ACC / AUC / AP) (%) on the OW-FFA-Bench and the test set of MA. * indicates the use of trained models provided by the authors, while the others were trained on the MA. Our method was trained on the FFA-VQA and the (image, answers) data from the fine-tuned MLLM, with all images sourced from MA. CLIPping uses the same visual encoder as our method, and we employ its best-performing approach (i.e., Prompt-tuning) for comparison. The bold and underline mark the best and second performances, respectively.
Experiments
In this section, we first establish the OW-FFA-Bench to evaluate the model’s generalization and robustness. We then compare our method with the state-of-the-art methods on this benchmark and conduct ablation studies. Furthermore, we perform qualitative studies by comparing FFAA with advanced MLLMs and employing several visualizations.
Experimental Setup
Datasets.
@@
We create the generalization test sets by randomly collecting face images from the public datasets: Deeperforensics (DPF) (Jiang et al. 2020), DeepFakeDetection (DFD) (Google 2019), DFDC (Dolhansky et al. 2020), PGGAN (Karras et al. 2017), Celeb-A (Liu et al. 2018), WhichFaceIsReal (WFIR) (West and Bergstrom 2019) and MultiFFDI (MFFDI) (on the Bund 2024). The images are then divided into six generalization test sets based on their source as shown in Tab. 1, with each containing 1K images and differing in distribution. See Appendix for more details.
Implementation details.
@@
We first fine-tune an MLLM, e.g. LLaVA-v1.6-mistral-7B (Liu et al. 2024a) with LoRA (rank=32, alpha=48) on the FFA-VQA dataset using 2 RTX 3090 GPUs for 3 epochs with a learning rate of 1e-4 and a batch size of 16. Then, we train MIDS using 2 RTX 3090 GPUs for 2 epochs with a learning rate of 1e-4 and a batch size of 48. See Appendix for more training details.
Evaluation metrics.
@@
We utilize the Accuracy (ACC), Area Under Receiver Operating Characteristic Curve (AUC) and Average Precision (AP) to empirically evaluate generalization performance on individual test sets as well as across the entire OW-FFA-Bench (ALL). Additionally, we employ the standard deviation of ACC (sACC) across all test sets to assess the model’s robustness.
Comparison with Existing Methods
In this section, we compare our method against state-of-the-art face forgery analysis methods on 1K in-domain test set of MA and the OW-FFA-Bench. These methods can be divided into two categories: methods using only visual information (i.e., Xception (Chollet 2017), RECCE (Cao et al. 2022), Implicit (Dong et al. 2023), FedForgery (Liu et al. 2023), Exposing (Ba et al. 2024) and NPR (Tan et al. 2024)), and methods based on CLIP using both textual and visual information (i.e., CLIPping (Khan and Dang-Nguyen 2024)).
Tab. 2 presents the comparison results. Methods in the first category display varying degrees of overfitting and poor generalization, underscoring the challenges in open-world scenarios where various confounding factors complicate the discrimination using solely visual information. Conversely, CLIPping effectively prevent overfitting and demonstrate better generalization. However, the accuracy and robustness of these methods remain suboptimal. In contrast, we frame face forgery analysis as a VQA task, requiring the model to analyze forgery before making a judgment. This approach improves the model’s generalization ability and clarifies its decision-making process. Given the fuzzy boundaries between real and forged faces, we employ hypothetical prompts and MIDS to comprehensively compare answers based on varying hypotheses, boosting robustness.
Figure 6: The distribution of ’Easy’ and ’Hard’ images as determined by the MLLM and the accuracy improvements on ’Hard’ images achieved by MIDS. On ’Easy’ images, performance is consistent both with and without MIDS.
Ablation Study
Ablation Study on FFA-VQA.
We study the effectiveness of the FFA-CoT and the impact of high-quality image descriptions and forgery reasoning textual data on various MLLMs. Specifically, we utilize advanced and open-source MLLMs, including LLaVA-phi-3-mini (Abdin et al. 2024) and LLaVA-mistral-7B (Liu et al. 2024a). We fine-tune these MLLMs on the FFA-VQA containing only ’Analysis results’ (R), ’Image description + Analysis results’ (DR), ’Forgery reasoning + Analysis results’ (FR) and the complete FFA-CoT (All) respectively. We evaluate the fine-tuned MLLMs on both the test set of MA and the OW-FFA-Bench. Fig. 3 presents the detection accuracy (%) of these MLLMs fine-tuned on corresponding datasets. Using only the classification result as the output of the MLLM yields overfitting and poor generalization. With the inclusion of image descriptions and forgery reasoning, the task evolves into a VQA task that necessitates deeper analysis and reasoning, enhancing generalization on unseen datasets. In addition, employing FFA-CoT further improves generalization.

### v2

@@ 4.2 Multi-answer Intelligent Decision System
MIDS consists of three components: Feature Encoding, Image-Answer Cross Fusion Module and Classification Layer, as depicted in Fig. 4b. We first utilize the fine-tuned MLLM to extract answers from unused samples in the Multi-attack dataset for training. Specifically, for a given image $\mathbf{X}_{v}$, a non-hypothetical prompt is randomly selected to query the MLLM, yielding $\mathbf{X}_{a}^{a}$. Based on the determination in $\mathbf{X}_{a}^{a}$, two hypothetical prompts assuming the same or opposite determination are used to query again, yielding the positive answer $\mathbf{X}_{a}^{p}$ and the negative answer $\mathbf{X}_{a}^{n}$ respectively. Finally, based on the authenticity $y_{v}\in\{0,1\}$ of $\mathbf{X}_{v}$ and the determination result $y_{a}\in\{0,1\}$ of $\mathbf{X}_{a}$, we assign a label $y$ for each $(\mathbf{X}_{v},\mathbf{X}_{a})$ as:
@@
Feature Encoding: For each given answer $\mathbf{X}_{a}$, to prevent the model from relying solely on the final classification result while ignoring the intermediate analysis process, we mask the ’Analysis result’ of $\mathbf{X}_{a}$. Then, we employ the pre-trained T5-Encoder [35] with frozen weights for good linguistic consistency to obtain the text feature $\mathbf{F}_{t}$. For the image $\mathbf{X}_{v}$, we employ the pre-trained CLIP-ViT-L/14 [34] as the visual feature encoder. To maintain basic image understanding and enhance the perception of local and global features related to face authenticity, the weights of the last two layers are unfrozen. Then, trainable projection matrices $(\mathbf{W}_{vg},\mathbf{W}_{vl})$ are employed to project the output features of the last two layers into the same dimensional space as $\mathbf{F}_{t}$, resulting in $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$.
@@
Image-Answer Cross Fusion: To extract the corresponding visual features based on the answer and to identify matching textual features from the visual features, we employ dual cross-attention for comprehensive fusion across all dimensions. Specifically, it can be expressed as:
$\mathbf{F}_{v\to t}=\text{LayerNorm}(\text{softmax}(\frac{\mathbf{Q}_{v}\mathbf{K}_{t}^{T}}{\sqrt{d}})\mathbf{V}_{t}),$ (4)
$\mathbf{F}_{t\to v}=\text{LayerNorm}(\text{softmax}(\frac{\mathbf{Q}_{t}\mathbf{K}_{v}^{T}}{\sqrt{d}})\mathbf{V}_{v}),$
@@
where $d$ is the dimension of these features, $\{\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\}=\mathcal{F}_{qkv}^{t}(\mathbf{F}_{t})$, $\{\mathbf{Q}_{v},\mathbf{K}_{v},\mathbf{V}_{v}\}=\mathcal{F}_{qkv}^{v}(\mathbf{F}_{v})$. We perform the operations of Eq. 4 with $(\mathbf{F}_{t},\mathbf{F}_{vl})$ and $(\mathbf{F}_{t},\mathbf{F}_{vg})$ respectively, to obtain $\mathbf{F}_{t\to vl}$, $\mathbf{F}_{vl\to t}$, $\mathbf{F}_{t\to vg}$ and $\mathbf{F}_{vg\to t}$. Subsequently, we employ learnable vectors $\mathbf{f}^{0}$ to obtain the corresponding aggregated feature based on cross-attention:
$\mathbf{f}^{r}=\text{LayerNorm}(\text{softmax}(\frac{\mathbf{Q}_{f}\mathbf{K}_{F}^{T}}{\sqrt{d}})\mathbf{V}_{F}),$ (5)
$\mathbf{Q}_{f}=\mathbf{W}_{f}^{Q}\mathbf{f}^{r-1},~{}\mathbf{K}_{F}=\mathbf{W}_{F}^{K}\mathbf{F},~{}\mathbf{V}_{F}=\mathbf{W}_{F}^{V}\mathbf{F},~{}$
@@
where $r$ denotes the $r$-th iteration. Through the operations of Eq. 5, we obtain text feature-enhanced visual domain features $(\mathbf{f}_{t\to vl}^{r},\mathbf{f}_{t\to vg}^{r})$ and image feature-enhanced textual domain features $(\mathbf{f}_{v\to tl}^{r},\mathbf{f}_{v\to tg}^{r})$ for each $(\mathbf{X}_{v},\mathbf{X}_{a})$.
@@
Classification and Loss Function: To select the best answer, the model must possess a certain ability to discern the authenticity of faces and answers. Therefore, we set the classification labels as the four-class categorization shown in Eq. 3. The [CLS] tokens of $\mathbf{F}_{vl}$ and $\mathbf{F}_{vg}$, containing enriched image category information, are passed through projection matrices $(\mathbf{W}_{vl}^{cls},\mathbf{W}_{vg}^{cls})$ to obtain $(\mathbf{f}_{vl}^{cls},\mathbf{f}_{vg}^{cls})$. Then, we concatenate $(\mathbf{f}_{vg}^{cls},\mathbf{f}_{t\to vg}^{r},\mathbf{f}_{v\to tg}^{r},\mathbf{f}_{vl}^{cls},\mathbf{f}_{t\to vl}^{r},\mathbf{f}_{v\to tl}^{r})$ to obtain $\mathbf{H}_{f}$, which are further enhanced by $M$ iterations of self-attention with residual connections [12], resulting in $\mathbf{H}^{\prime}_{f}$. Finally, $\mathbf{H}^{\prime}_{f}$ is passed through $\mathbf{W}_{c}$ and a softmax to obtain $\mathbf{m}\in\mathbb{R}^{4}$. The loss is computed as follows:
@@
[…]
Naive ResNet-34[12] 33.8 / 30.2 74.6 / 80.3 58.0 / 69.6 36.1 / 55.4 78.1 / 92.2 64.1 / 70.8 57.4 / 61.1 95.6 / 99.2 20.8
Naive EfficientNet-B4[42] 30.3 / 24.8 79.8 / 87.2 51.5 / 75.0 62.5 / 94.9 96.5 / 99.5 78.0 / 86.7 66.4 / 75.2 98.8 / 99.9 22.9
Naive Xception[6] 61.1 / 60.2 78.0 / 87.8 45.7 / 74.8 69.4 / 96.5 91.7 / 98.5 79.3 / 86.5 70.9 / 75.2 96.8 / 99.4 16.3
Spatial RECCE[4] 39.9 / 28.7 72.2 / 82.8 41.5 / 73.7 93.6 / 98.0 91.6 / 97.4 75.9 / 83.6 69.1 / 74.0 98.1 / 99.6 22.4
Spatial DCL[39] 57.9 / 63.9 82.5 / 89.3 47.5 / 78.5 41.2 / 99.2 98.0 / 99.8 78.1 / 89.8 67.5 / 76.2 98.2 / 99.6 21.6
Spatial Implicit*[9] 46.7 / 46.5 77.5 / 87.7 53.7 / 70.8 56.6 / 67.0 49.6 / 51.1 67.8 / 78.1 58.7 / 63.3 67.8 / 70.2 10.4
Spatial FedForgery*[25] 75.8 / 83.5 76.6 / 88.1 38.6 / 58.2 68.5 / 55.9 48.9 / 47.8 52.5 / 56.8 60.1 / 67.8 51.0 / 64.6 13.7
Spatial Exposing[3] 64.2 / 75.5 76.4 / 86.2 48.6 / 77.2 62.3 / 94.8 90.0 / 98.0 75.2 / 83.4 69.5 / 76.8 94.5 / 98.8 14.9
Spatial NPR[41] 47.5 / 44.2 67.2 / 74.5 40.1 / 66.4 67.2 / 80.1 91.6 / 99.3 71.4 / 79.4 64.2 / 72.5 97.1 / 99.7 19.3
Spatial UIA-ViT[50] 69.3 / 74.6 80.8 / 88.5 55.6 / 87.6 79.9 / 99.4 95.1 / 99.1 83.1 / 90.7 77.3 / 84.4 98.4 / 99.8 13.6
Spatial CLIPping[18] 90.4 / 97.6 77.8 / 86.6 64.0 / 71.9 46.0 / 97.1 91.5 / 98.0 55.6 / 96.4 70.9 / 80.9 87.4 / 94.2 16.9
Frequency F3Net[33] 52.2 / 52.7 77.9 / 84.6 58.8 / 79.9 75.1 / 96.5 87.4 / 95.1 80.9 / 89.0 72.1 / 80.1 96.8 / 99.5 14.4
Frequency Two-Stream[29] 50.8 / 57.7 75.0 / 80.6 47.1 / 72.1 52.7 / 87.9 86.4 / 96.4 77.8 / 85.5 65.0 / 72.5 97.5 / 99.6 18.1
Frequency FADD[40] 59.0 / 61.6 67.6 / 69.6 76.3 / 59.4 14.8 / 20.0 86.3 / 97.7 62.5 / 70.7 61.1 / 68.9 89.6 / 94.0 23.2
Ours FFAA (w/o MIDS) 95.0 / 96.0 81.6 / 83.9 53.6 / 72.4 54.7 / 75.2 98.4 / 98.8 67.7 / 73.7 75.2 / 77.8 87.0 / 88.7 17.0
Ours FFAA 94.2 / 98.8 82.7 / 92.0 65.0 / 74.0 91.8 / 99.7 98.2 / 99.8 87.1 / 98.6 86.5 / 94.4 86.8 / 96.7 10.0
Table 1: Comparison results (ACC / AUC) (%) on OW-FFA-Bench and the test set of MA. * denotes the use of trained models provided by the authors, while the others were trained on MA. Our method was trained on FFA-VQA and (image, answers) data from the fine-tuned MLLM, with all images sourced from MA. Comparing our final method with other methods, bold and underline highlight the best and second-best performances, respectively. The gray-shaded areas indicate the performance of FFAA without MIDS.
5 Experiments
In this section, we first establish OW-FFA-Bench to evaluate model generalization and robustness. We then compare our method with the state-of-the-art methods on this benchmark and conduct ablation studies. Furthermore, we perform qualitative studies by comparing FFAA with advanced MLLMs and employing several visualizations.
@@ 5 Experiments
Source Deeperforensics[15] DeepFakeDetection[11] DFDC[8]
@@
Source PGGAN[16], Celeb-A[28] WhichFaceIsReal[46] MultiFFDI[31]
@@
5.1 Experimental Setup
@@
Datasets: We create the generalization test sets by randomly collecting face images from seven public datasets. The images are then divided into six test sets based on their source as shown in Tab. 2, with each containing 1K images and differing in distribution.
@@
Implementation details: We first fine-tune an MLLM, e.g. LLaVA-v1.6-mistral-7B [26] with LoRA (rank=32, alpha=48) on FFA-VQA using 2 RTX 3090 GPUs for 3 epochs with a learning rate of 1e-4 and a batch size of 16. Then, we train MIDS using 2 RTX 3090 GPUs for 2 epochs with a learning rate of 1e-4 and a batch size of 48. See Supplementary Material for more training details.
@@
Evaluation metrics: We utilize Accuracy (ACC) and Area Under Receiver Operating Characteristic Curve (AUC) to empirically evaluate generalization performance on individual test sets and across the entire OW-FFA-Bench (ALL). Additionally, we employ the standard deviation of ACC (sACC) across all test sets to assess model robustness.
Figure 5: Qualitative examples in real-world scenarios. Left: a facial frame from a video with a deepfake face of Zelenskyy delivering false statements. Right: a facial frame from a video of Zelenskyy’s genuine speech. Judgment results are indicated in parentheses as (’Real’, ’Fake’, ’Refuse to judge’), with green for correct judgments and red for incorrect ones.
5.2 Comparison with Competitive Methods
In this section, we compare our method against state-of-the-art face forgery analysis methods on 1K in-domain test set of MA and OW-FFA-Bench. These methods fall into three categories: (1) Naive: utilizing conventional classification models for binary classification; (2) Spatial: incorporating specialized modules or additional constraints to facilitate the extraction of forgery cues at the spatial level, followed by classification; (3) Frequency: leveraging the frequency domain to identify forgery characteristics, which are then employed for classification.
Tab. 1 presents the comparison results. These three categories of methods exhibit varying degrees of overfitting and limited generalization (i.e., strong performance on in-domain MA yet significantly reduced on cross-domain OW-FFA-Bench). Additionally, substantial variability in detection accuracy across different test sets reveals limited model robustness (i.e., demonstrating high sACC). In such cases, providing only binary classification results poses significant risks in practical applications. In contrast, we frame face forgery analysis as a VQA task, requiring forgery analysis prior to judgment, which improves generalization and clarifies the decision-making process. To address fuzzy boundaries between real and forged faces, we integrate hypothetical prompts and MIDS for a comprehensive comparison of answers under different hypotheses, enhancing robustness. Overall, our method significantly improves accuracy and robustness while clarifying the model’s decision-making, enabling more effective application in real-world scenarios.